Find the equation( in slope intercept form)of the line passing through the points given coordinates(-4,0) ,(-1,2)
Answer by mananth(16949)

slope = y2-y1/x2-x1
=2-0/-1-(-4)
=2/3
...
y=mx+b
Take any point (-4,0)
plug x,y,m to get b
0=2/3*-4 +b
0=-8/3
b= 8/3
...
equation = y=(2x/3)+8/3
...
m.ananth@hotmail.ca